OpenStudy (anonymous):

Translate the sentence into an equation. Four more than a number is 12. A. x + 12 = 4 B. x – 4 = 12x C. x – 4 = 12 + x D. x + 4 = 12 @sambolin1996

OpenStudy (anonymous):

D.

OpenStudy (anonymous):

d

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

yep what

OpenStudy (anonymous):

can someone medal me

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you got it right

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

Okay, this is a hard one

OpenStudy (anonymous):

how did you get two medels

OpenStudy (anonymous):

how hard

OpenStudy (anonymous):

A

OpenStudy (anonymous):

easy in a way

OpenStudy (anonymous):

im not sure if im right

OpenStudy (anonymous):

that one is hard

OpenStudy (anonymous):

Even for me

OpenStudy (anonymous):

No like you ahve to amtch them up

OpenStudy (anonymous):

hah its fine moving on

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Heres another one

OpenStudy (anonymous):

D or either A

OpenStudy (anonymous):

But im going more on D

OpenStudy (anonymous):

Choose the equation that models the problem. Zach's baseball team won 32 games last season. They played a total of 45 games. How many games did they lose overall? Let l represent the number of games that Zach's team lost. A. l – 32 = 45 B. l + 32 = 45 C. l + 45 = 32 D. 32l = 45

OpenStudy (anonymous):

all those question marks represent a

OpenStudy (anonymous):

i meant a -

OpenStudy (anonymous):

a

OpenStudy (anonymous):

kk 1 more

OpenStudy (anonymous):

Charlie has 8 bills in his wallet. Five of the bills are $10 bills, and the other bills are $5 bills. What is the total value of the 8 bills in Charlie's wallet? The total value of the 8 bills is $

OpenStudy (anonymous):

what is the totoal of $?

OpenStudy (anonymous):

@KailaMorgan7777

OpenStudy (anonymous):

$65

OpenStudy (anonymous):

for what